# EzSQL: An SQL intermediate representation for improving SQL-to-text Generation

### v1

@@ Abstract
The SQL-to-text generation task traditionally uses template base, Seq2Seq, tree-to-sequence, and graph-to-sequence models. Recent models take advantage of pre-trained generative language models for this task in the Seq2Seq framework. However, treating SQL as a sequence of inputs to the pre-trained models is not optimal. In this work, we put forward a new SQL intermediate representation called EzSQL to align SQL with the natural language text sequence. EzSQL simplifies the SQL queries and brings them closer to natural language text by modifying operators and keywords, which can usually be described in natural language. EzSQL also removes the need for set operators. Our proposed SQL-to-text generation model uses EzSQL as the input to a pre-trained generative language model for generating the text descriptions. We demonstrate that our model is an effective state-of-the-art method to generate text narrations from SQL queries on the WikiSQL and Spider datasets. We also show that by generating pretraining data using our SQL-to-text generation model, we can enhance the performance of Text-to-SQL parsers.
@@ 1 Introduction
An SQL-to-text model automatically generates human-like text descriptions by interpreting the meaning signified by the structured query language (SQL) query. This task has been crucial for the natural language to interact with a database as it helps users who are unfamiliar with SQL to understand the queries. A newly found application of this task is training data generation for enhancing the performance of Text-to-SQL parsers [17, 20]. Earlier attempts were made for the SQL-to-text task in the form of rule-based and template-based [10, 14] models. Further, newer approaches use sequence-to-sequence (Seq2Seq) networks to model the SQL queries and natural language in an interconnected manner [7]. Furthermore, several models were also developed based on graph encoding techniques [9, 18], since SQL is designed to express the query intent in a graph-structured manner.
Recently, self-supervised deep contextual language models [2, 13, 11] have shown their effective modeling ability for text, demonstrating state-of-the-art results in a series of NLP tasks. These tasks include sequence understanding and language generation. Previously, [17, 20] have explored the use of BART [11] in SQL-to-Text generation. To generate the text utterances, the authors train their SQL-to-Text model where the input is the original SQL. The input SQL is directly taken as a sequence by the BART tokenizer since the pre-trained model expects a sequence of text as input. However, directly feeding SQL as a sequence is not optimal.
@@
In this paper, we present EzSQL, a new intermediate representation of queries focusing on the SQL-to-text task. EzSQL bridges the gap between SQL and natural language (NL). Figure 1 presents the EzSQL-based SQL-to-text translation. In the example, EzSQL uses the OR conjunctive operator instead of the UNION operator, which also removes the multiple SELECT clauses, and aligns better with the natural language. These simplification steps in EzSQL make it more suitable as an input to a pre-trained language model for the generation of the text descriptions. We propose an SQL-to-Text generator that leverages the capability of a pre-trained language model (BART) to generate text from a given SQL query. We first translate the SQL queries into EzSQL which is more suitable for the pre-trained language model. The EzSQL is fed as an input to BART, where it is directly tokenized by the BART tokenizer without additional pre-processing and the BART decoder generates the text descriptions.
@@ 2 Related Work
Data augmentation for semantic parsing has attracted a lot of attention in recent years. Earlier attempts include generating new examples using SCFG [8], back-translation [16, 3] etc. [27, 25] also use handcrafted SCFG rules to generate new parallel data. In another approach, [29] attempted to generate a large number of SQL queries based on an abstract syntax tree grammar and used a hierarchical SQL-to-text generation model in order to synthesize natural language questions. However, this approach uses the traditional seq2seq mechanism for the question generation. Recent approaches such as [17, 20] use SQL-to-text for parallel data generation by using powerful conditional text generators such as BART. Following these recent approaches, we also utilize BART for SQL-to-text and data augmentation.
@@
Significant progress has been achieved in text-to-SQL over the past few years. With a suitable amount of in-domain training data, existing models already achieve over 80% exact matching accuracy [21] on single-domain benchmarks like ATIS and GeoQuery. However, annotating NL questions with SQL queries is resource intensive which makes it cost-inefficient to collect training examples from all possible databases. [28] introduced Spider, a cross-database text-to-SQL benchmark evaluates a system using different databases. Recently, several models that target the Spider benchmark have been introduced, including RAT-SQL [19] and several other models that build on top of RAT-SQL [17, 4, 20].
@@ 3.1 Overview
(5) Uses DISTINCT and COUNT keywords in a manner that inclines with the natural language.
@@ 3.2 EzSQL simplification procedure
The simplification process in EzSQL can comprise of several steps depending on the type and complexity of the query. EzSQL takes several inspirations from existing IR used for Text-to-SQL. The IR in [26] represents SQL statements without FROM and JOIN ON clauses. [5] removes the FROM, JOIN ON and GROUP BY clauses, and combines the WHERE and HAVING conditions. The recent IR in [4] dispenses nested subqueries and set operators. By taking advantage of the existing IRs and combining the generation processes, we implement the EzSQL IR for SQL-to-text with the single goal of representing the SQL statement closest to natural language text. Since our representation does not have any limitations and restrictions for simplification, the implementation is much simpler in the one-way process. The simplification process is represented in figure 5.
@@ 12       Set nested_query_flag equal to 2
These split queries are free of set operators. Since they are independent in nature, it makes them simpler to process, while at the same time inclines them towards natural language. After this, further processing takes place on these split queries independently.
@@ 3       Find the occurrence of each user-defined name;
4       Replace each occurrence of user-defined name, with the corresponding original name
@@
Algorithm 4 makes the use of the lists generated by algorithm 3. It finds the occurrence of each user-defined name( i.e. alias) and replaces it with the corresponding original table name, by referring to the original_names list. Thus, all aliases are replaced by their respective original table names.
@@ 5       Output queries are input for the next stage
Algorithm 5 is the final step associated with the processing of aliases. The output strings of algorithm 4 are the input for this algorithm. These input strings contain the ‘as’ keyword, and the original table name occurs before and after it ( because of Algorithm 4) Thus, the ‘as’ keyword is identified, and is removed, along with the word occurring after it. Thus, we obtain queries without aliases and ‘as’ keywords.
@@ 3.3 SQL-to-text Generation Model
Our SQL-to-text generation approach takes an SQL query as the input and outputs a natural language question. Our approach is based on pre-trained sequence-to-sequence BART model [11]. The model consists of an SQL simplifier module to generate EzSQL and a BART encoder-decoder module. The SQL simplifier takes SQL as an input and produces the EzSQL IR. The EzSQL IR is then taken by the BART model to generate text descriptions.
@@ 4.1 SQL-to-text
We evaluate our model on two datasets, WikiSQL [31] and Spider [28]. WikiSQL consists of a corpus of 87,726 hand-annotated SQL queries and natural language question pairs. These SQL queries are further split into training (61,297 examples), development (9,145 examples), and test sets (17,284 examples). Spider consists of 7000, 1034, and 2147 samples for training, development, and testing respectively, where 206 databases are split into 146 for training, 20 for development, and 40 for testing. We use the BART-large variant in our model. The BLEU-4 score [15] is used as the automatic evaluation metric for this task.
@@
Results and Discussion.
Table 1 summarizes the results of our models and baselines. On the WikiSQL dataset, our model outperforms the existing approaches that use different kinds of embedding methods. This shows that the EzSQL IR can bypass the need for any special embedding methods such as Tree or Graph embeddings for the SQL. Although several SQL-to-text models exist for the simple WikiSQL dataset, few models exist for the smaller and more complex Spider dataset. Recent, approaches use BART as the translator, mainly because generating texts using such small a dataset is not trivial, and taking advantage of pre-trained models can be seen as one of the solutions. The results show that our model performs significantly better than the existing baseline on this dataset. Both models use the same BART-large model, however we use the EzSQL IR as the input instead of the raw SQL. Our approach to bridging the gap between SQL and the NL targets using EzSQL is important as the pre-trained encoders expect a string of text as input whereas the existing approach of feeding complex SQLs as a string of text further complicates the task of the encoder. Our model pipeline with EzSQL is much simpler and easier to train since it can efficiently reuse all the parameters of the pre-trained Seq2seq model and does not introduce any new model parameters.
@@
Model Error Analysis.
We analyze the probable causes for error in text generation with BERT sentence similarity score using Sentence BERT. We compute the sentence similarity scores between the gold sentence and the generated sentence. For comparison, we also compute the similarity between the gold and other manually corrected as well as incorrect narrations of the SQL. The comparison of the scores is shown in Figure 7. The erroneous narration and the corrected version have almost the same similarity score. This indicates that our model will also not have enough information to correct the error based on the target only. In the last two examples, a few alterations in the narration can change or flip the scores, confusing the model and leading to the incorrect version as output. Here, the input query is “SELECT count(*) FROM FLIGHTS AS T1 JOIN AIRPORTS AS T2 ON T1.DestAirport = T2.AirportCode JOIN AIRPORTS AS T3 ON T1.SourceAirport = T3.AirportCode WHERE T2.City = “Ashley” AND T3.City = “Aberdeen””, where the order of information is “to-from”. However, we expect the language models to have an overwhelming majority of examples in the “from-to’ order. Hence, the generated text is in the incorrect order and the target loss cannot correct the error as we have seen in the analysis. The only way to correct the generation is by reordering the information according to the bias of language models in the intermediate representation.
@@ 4.2 Text-to-SQL
We further experiment on the effectiveness of our model in pretraining data generation for Text-to-SQL parsers. Following [20], we use the database-specific probabilistic context-free grammar (PCFG) model to sample new SQL queries from the Spider dataset. In this method, ASDL [22] grammar of SQL is represented by a set of context-free grammar (CFG) rules. The grammar is pre-defined based on expert knowledge and it has been successfully applied in previous work [24, 19] on Text-to-SQL parsing. A separate PCFG is trained on each source database based on the assumption that each database has a different distribution of SQL queries. Using our SQL-to-text model trained on the Spider dataset, we generate the questions from the newly sampled SQLs.
@@
Results and Discussion.
The evaluation results of the parser are shown in Table 2. It can be seen that the RAT-SQL parser pre-trained using the synthesized data with the help of our SQL-to-text model improves the performance of the RAT-SQL baseline. Moreover, the parser outperforms the pre-trained RAT-SQL of [20] although the same method of data synthesis is used in both models.
The results in Table 2 also serve as an ablation study. Here, we identify whether the EzSQL IR method is necessary and does it helps in improving the final Text-to-SQL parsing performance. Although in the SQL-to-text experiments, we have seen significant improvement in performance, the goal of the SQL-to-text task, going forward, is assisting in data augmentation and data generation for semantic parsers. The improvement of performance over the existing RAT-SQL pre-train model by introducing EzSQL IR as the only additional factor shows the effectiveness of our enhanced SQL-to-text model in data generation.
@@ 5 Conclusion and Future Work
We proposed a new SQL intermediate representation aimed to improve the SQL-to-text task that is simple yet effective for enhancing translation performance using pre-trained language models. The proposed SQL-to-text model was able to outperform the baselines on both WikiSQL and Spider datasets. Moreover, we successfully show that the proposed model can be used to synthesize pre-training data to further enhance the performance of Text-to-SQL semantic parsers. In the future, we also intend on adding natural keywords to EzSQL, which relay the semantic sense more clearly, that current LLMs tend to struggle with. For example, when given the SQL -“ SELECT T1.AirportCode FROM AIRPORTS AS T1 JOIN FLIGHTS AS T2 ON T1.AirportCode = T2.DestAirport OR T1.AirportCode = T2.SourceAirport GROUP BY T1.AirportCode ORDER BY count(*) LIMIT 1 ” to GPT-4, it gives the output as - “Which airport is involved in the highest number of flights, either as a source or destination airport?”. This is wrong, as the order by is ascending by default, and the output should be the fewest number of flights. Thus, these errors would be avoided if a natural language keyword (like ascending in this case) is included in an intermediate form, like EzSQL.

### v2

@@ Abstract
The SQL-to-text generation task traditionally uses template base, Seq2Seq, tree-to-sequence, and graph-to-sequence models. Recent models take advantage of pre-trained generative language models for this task in the Seq2Seq framework. However, treating SQL as a sequence of inputs to the pre-trained models is not optimal. In this work, we put forward a new SQL intermediate representation called EzSQL to align SQL with the natural language text sequence. EzSQL simplifies SQL queries and brings them closer to natural language text by modifying operators and keywords, which can usually be described in natural language. EzSQL also removes the need for set operators. Our proposed SQL-to-text generation model uses EzSQL as the input to a pre-trained generative language model for generating the text descriptions. We demonstrate that our model is an effective state-of-the-art method to generate text narrations from SQL queries on the WikiSQL and Spider datasets. We also show that by generating pretraining data using our SQL-to-text generation model, we can enhance the performance of Text-to-SQL parsers.
@@ 1 Introduction
An SQL-to-text model automatically generates human-like text descriptions by interpreting the meaning signified by the structured query language (SQL) query. This task has been crucial for natural language to interact with a database, as it helps users who are unfamiliar with SQL to understand the queries. A newly found application of this task is training data generation for enhancing the performance of Text-to-SQL parsers [17, 20]. Earlier attempts were made for the SQL-to-text task in the form of rule-based and template-based [10, 14] models. Further, newer approaches use sequence-to-sequence (Seq2Seq) networks to model the SQL queries and natural language in an interconnected manner [7]. Furthermore, several models were also developed based on graph encoding techniques [9, 18], since SQL is designed to express the query intent in a graph-structured manner.
Recently, self-supervised deep contextual language models [2, 13, 11] have shown their effective modeling ability for text, demonstrating state-of-the-art results in a series of NLP tasks. These tasks include sequence understanding and language generation. Previously, [17, 20] have explored the use of BART [11] in SQL-to-Text generation. To generate the text utterances, the authors train their SQL-to-Text model, where the input is the original SQL. The input SQL is directly taken as a sequence by the BART tokenizer since the pre-trained model expects a sequence of text as input. However, directly feeding SQL as a sequence is not optimal.
@@
In this paper, we present EzSQL, a new intermediate representation of queries focusing on the SQL-to-text task. EzSQL bridges the gap between SQL and natural language (NL). Figure 1 presents the EzSQL-based SQL-to-text translation. In the example, EzSQL uses the OR conjunctive operator instead of the UNION operator, which also removes the multiple SELECT clauses, and aligns better with the natural language. These simplification steps in EzSQL make it more suitable as an input to a pre-trained language model for the generation of text descriptions. We propose an SQL-to-Text generator that leverages the capability of a pre-trained language model (BART) to generate text from a given SQL query. We first translate the SQL queries into EzSQL, which is more suitable for the pre-trained language model. The EzSQL is fed as input to BART, where it is directly tokenized by the BART tokenizer without additional pre-processing, and the BART decoder generates the text descriptions.
@@ 2 Related Work
Data augmentation for semantic parsing has attracted a lot of attention in recent years. Earlier attempts include generating new examples using SCFG [8], back-translation [16, 3], etc. [27, 25] also use handcrafted SCFG rules to generate new parallel data. In another approach, [29] attempted to generate a large number of SQL queries based on an abstract syntax tree grammar and used a hierarchical SQL-to-text generation model in order to synthesize natural language questions. However, this approach uses the traditional seq2seq mechanism for the question generation. Recent approaches, such as [17, 20] use SQL-to-text for parallel data generation by using powerful conditional text generators such as BART. Following these recent approaches, we also utilize BART for SQL-to-text and data augmentation.
@@
Significant progress has been achieved in text-to-SQL over the past few years. With a suitable amount of in-domain training data, existing models already achieve over 80% exact matching accuracy [21] on single-domain benchmarks like ATIS and GeoQuery. However, annotating NL questions with SQL queries is resource-intensive, which makes it cost-inefficient to collect training examples from all possible databases. [28] introduced Spider, a cross-database text-to-SQL benchmark evaluates a system using different databases. Recently, several models that target the Spider benchmark have been introduced, including RAT-SQL [19] and several other models that build on top of RAT-SQL [17, 4, 20].
@@ 3.1 Overview
(5) Uses DISTINCT and COUNT keywords in a manner that aligns with the natural language.
@@ 3.2 EzSQL simplification procedure
The simplification process in EzSQL can comprise several steps depending on the type and complexity of the query. EzSQL takes several inspirations from existing IR used for Text-to-SQL. The IR in [26] represents SQL statements without FROM and JOIN ON clauses. [5] removes the FROM, JOIN ON and GROUP BY clauses, and combines the WHERE and HAVING conditions. The recent IR in [4] dispenses nested subqueries and set operators. By taking advantage of the existing IRs and combining the generation processes, we implement the EzSQL IR for SQL-to-text with the single goal of representing the SQL statement closest to natural language text. Since our representation does not have any limitations and restrictions for simplification, the implementation is much simpler in the one-way process. The simplification process is represented in figure 5.
@@ 12       Set nested_query_flag equal to 2
These split queries are free of set operators. Since they are independent in nature, it makes them simpler to process, while at the same time inclining them towards natural language. After this, further processing takes place on these split queries independently.
@@ 3       Find the occurrence of each user-defined name;
4       Replace each occurrence of the user-defined name with the corresponding original name
@@
Algorithm 4 makes the use of the lists generated by algorithm 3. It finds the occurrence of each user-defined name (i.e., alias) and replaces it with the corresponding original table name, by referring to the original_names list. Thus, all aliases are replaced by their respective original table names.
@@ 5       Output queries are input for the next stage
Algorithm 5 is the final step associated with the processing of aliases. The output strings of algorithm 4 are the input for this algorithm. These input strings contain the ‘as’ keyword, and the original table name occurs before and after it ( because of Algorithm 4). Thus, the ‘as’ keyword is identified, and is removed, along with the word occurring after it. Thus, we obtain queries without aliases and ‘as’ keywords.
@@ 3.3 SQL-to-text Generation Model
Our SQL-to-text generation approach takes an SQL query as input and outputs a natural language question. Our approach is based on a pre-trained sequence-to-sequence BART model [11]. The model consists of an SQL simplifier module to generate EzSQL and a BART encoder-decoder module. The SQL simplifier takes SQL as input and produces the EzSQL IR. The EzSQL IR is then taken by the BART model to generate text descriptions.
@@ 4.1 SQL-to-text
We evaluate our model on two datasets, WikiSQL [31] and Spider [28]. WikiSQL consists of a corpus of 87,726 hand-annotated SQL queries and natural language question pairs. These SQL queries are further split into training (61,297 examples), development (9,145 examples), and test sets (17,284 examples). Spider consists of 7000, 1034, and 2147 samples for training, development, and testing, respectively, where 206 databases are split into 146 for training, 20 for development, and 40 for testing. We use the BART-large variant in our model. The BLEU-4 score [15] is used as the automatic evaluation metric for this task.
@@
4.1.1 Results and Discussion
Table 1 summarizes the results of our models and baselines. On the WikiSQL dataset, our model outperforms the existing approaches that use different kinds of embedding methods. This shows that the EzSQL IR can bypass the need for any special embedding methods, such as Tree or Graph embeddings for the SQL. Although several SQL-to-text models exist for the simple WikiSQL dataset, few models exist for the smaller and more complex Spider dataset. Recently, approaches use BART as the translator, mainly because generating texts using such a small dataset is not trivial, and taking advantage of pre-trained models can be seen as one of the solutions. The results show that our model performs significantly better than the existing baseline on this dataset. Both models use the same BART-large model, however we use the EzSQL IR as the input instead of the raw SQL. Our approach to bridging the gap between SQL and the NL targets using EzSQL is important, as the pre-trained encoders expect a string of text as input, whereas the existing approach of feeding complex SQLs as a string of text further complicates the task of the encoder. Our model pipeline with EzSQL is much simpler and easier to train since it can efficiently reuse all the parameters of the pre-trained Seq2seq model and does not introduce any new model parameters.
@@ 4.1.1 Results and Discussion
4.1.2 Model Error Analysis
We analyze the probable causes for error in text generation with BERT sentence similarity score using Sentence BERT. We compute the sentence similarity scores between the gold sentence and the generated sentence. For comparison, we also compute the similarity between the gold and other manually corrected, as well as incorrect, narrations of the SQL. The comparison of the scores is shown in Figure 7. The erroneous narration and the corrected version have almost the same similarity score. This indicates that our model will also not have enough information to correct the error based on the target only. In the last two examples, a few alterations in the narration can change or flip the scores, confusing the model and leading to the incorrect version as output. Here, the input query is “SELECT count(*) FROM FLIGHTS AS T1 JOIN AIRPORTS AS T2 ON T1.DestAirport = T2.AirportCode JOIN AIRPORTS AS T3 ON T1.SourceAirport = T3.AirportCode WHERE T2.City = “Ashley” AND T3.City = “Aberdeen””, where the order of information is “to-from”. However, we expect the language models to have an overwhelming majority of examples in the “from-to’ order. Hence, the generated text is in the incorrect order, and the target loss cannot correct the error, as we have seen in the analysis. The only way to correct the generation is by reordering the information according to the bias of language models in the intermediate representation.
@@ 4.2 Text-to-SQL
We further experimented on the effectiveness of our model in pretraining data generation for Text-to-SQL parsers. Following [20], we use the database-specific probabilistic context-free grammar (PCFG) model to sample new SQL queries from the Spider dataset. In this method, ASDL [22] grammar of SQL is represented by a set of context-free grammar (CFG) rules. The grammar is pre-defined based on expert knowledge, and it has been successfully applied in previous work [24, 19] on Text-to-SQL parsing. A separate PCFG is trained on each source database based on the assumption that each database has a different distribution of SQL queries. Using our SQL-to-text model trained on the Spider dataset, we generate the questions from the newly sampled SQLs.
@@
4.2.1 Results and Discussion
The evaluation results of the parser are shown in Table 2. It can be seen that the RAT-SQL parser pre-trained using the synthesized data with the help of our SQL-to-text model improves the performance of the RAT-SQL baseline. Moreover, the parser outperforms the pre-trained RAT-SQL of [20], although the same method of data synthesis is used in both models.
The results in Table 2 identify whether the EzSQL IR method is necessary and if it helps to improve the final Text-to-SQL parsing performance. Although in the SQL-to-text experiments we have seen significant improvement in performance, the goal of the SQL-to-text task, going forward, is to assist in data augmentation and data generation for semantic parsers. The improvement of performance over the existing RAT-SQL pre-train model by introducing EzSQL IR as the only additional factor shows the effectiveness of our enhanced SQL-to-text model in data generation.
4.3 Ablation Study
Figure 8: Ablation study for SQL-to-Text translation task.
Figure 9: Ablation study for Text-to-SQL translation task.
We perform ablation studies for both the SQL-to-Text and Text-to-SQL tasks to understand the effect of our proposed method on the performance of the translations. Figure 8 shows the ablation study for SQL-to-text translation. We perform this study by taking the EzSQL model with BART and then analyzing the performance without the proposed EzSQL model. This is done for two standard datasets, viz. WikiSQL and Spider. BART was used in this experiment to make a fair comparison of the performance with existing models such as [11, 17] that use the same BART model.
Figure 9 shows the ablation study for Text-to-SQL translation. Similar to the previous study, we take the well-known RAT-SQL model pre-trained with our EzSQL IR and compare the performance without our method on the Spider dataset. In this experiment, we use the BERT and Electra models to match the settings of existing baseline and state-of-the-art models such as [20] and [29].
As illustrated in Figures 8 and 9, we see that the proposed model has a significant contribution to improving the performance of the SQL-to-text and Text-to-SQL translation.
@@ 5 Conclusion and Future Work
We proposed a new SQL intermediate representation aimed at improving the SQL-to-text task that is simple yet effective for enhancing translation performance using pre-trained language models. The proposed SQL-to-text model was able to outperform the baselines on both WikiSQL and Spider datasets. Moreover, we successfully show that the proposed model can be used to synthesize pre-training data to further enhance the performance of Text-to-SQL semantic parsers. The proposed EzSQL method also illustrates several advantages. EzSQL produces an intermediate representation (IR) of any SQL, and this IR can be used to train any existing seq2seq model and enhance its performance. The method is simple and has little overhead in the generation of the IRs. The flexibility and modularity introduced by the proposed method can enhance Text-to-SQL tasks as well by generating augmented data as shown in the experiments. However, EzSQL introduces an extra step in the end-to-end training of a seq2seq model, where we need to introduce the IR to the model being trained. Although the EzSQL model works for most of the popular Text-to-SQL datasets, some custom databases may need some additional settings in the proposed EzSQL model to produce the best results. In the future, we intend to add natural keywords to EzSQL, which relay the semantic sense more clearly, that current LLMs tend to struggle with. For example, when given the SQL -“ SELECT T1.AirportCode FROM AIRPORTS AS T1 JOIN FLIGHTS AS T2 ON T1.AirportCode = T2.DestAirport OR T1.AirportCode = T2.SourceAirport GROUP BY T1.AirportCode ORDER BY count(*) LIMIT 1 ” to GPT-4, it gives the output as - “Which airport is involved in the highest number of flights, either as a source or destination airport?”. This is wrong, as the order by is ascending by default, and the output should be the fewest number of flights. Thus, these errors would be avoided if a natural language keyword (like ascending in this case) is included in an intermediate form, like EzSQL.